OpenStudy (anonymous):

Total number of integral points lying inside the triangle formed by the lines \(xy=0\) and \(x+y=a\), where \(a\) is any positive integer.

OpenStudy (anonymous):

i just wanna clear my doubts for example xy=0 and x+y=3 one integral point inside this one ??

OpenStudy (anonymous):

Yeah, I think so but I couldn't be sure either I got it on my test today.

OpenStudy (anonymous):



OpenStudy (anonymous):

a=3 1 points a=4 1+2 points a=5 1+2+3 points ...

OpenStudy (anonymous):

For a = 21, 1 + 2 + 3 + ... + 19 = 19*20/2 = 190?

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

I happened to get the same but wasn't sure of it.